# Do MLLMs Capture How Interfaces Guide User Behavior? A Benchmark for Multimodal UI/UX Design Understanding

### v3

@@ 5 Experiments
We evaluate a diverse set of MLLMs, spanning both proprietary and open-source models. On the proprietary side, we include o1 (OpenAI et al. 2024b), GPT-4o (OpenAI et al. 2024a), Claude 3.5 Sonnet (Anthropic 2024), all noted for their strong visual reasoning. For open-source models, we evaluate widely adopted models such as Qwen2.5-VL (7B, 32B) (Bai et al. 2025), InternVL-2.5 (8B, 38B) (Chen et al. 2025), LLaVA-NeXT-7B (Liu et al. 2024a) and LLaVA-OneVision-7B (Li et al. 2024b).
@@ 5.1 UI/UX Design Selection
Model accuracies are reported across input-order variations of each UI image pair, reflecting the multi-image input setting. Specifically, we report: (1) First Accuracy (FA), when the more effective UI/UX design is presented first in the input; (2) Second Accuracy (SA), when it is presented second; and (3) Average Accuracy (AA), the mean of FA and SA, reflecting overall correctness. Finally, we compute (4) Consistent Accuracy (CA), which checks whether the model selects the same, correct UI across both input orders of each pair. This metric isolates the model’s core visual reasoning ability by ensuring that its choices are based on content rather than position bias, a well-known issue in pairwise comparison settings (Wang et al. 2025b; Zhang et al. 2023). (Random guessing yields a CA of 25%: since there’s a 50% chance of being correct in each order.) All evaluations are conducted over three independent runs, and we report the averaged results.
Model FA SA AA CA
Random 50.00 50.00 50.00 25.00
o1 16.56 97.78 57.17 15.56
GPT-4o 31.89 88.33 60.11 30.11
Claude 3.5 Sonnet 26.11 86.67 56.39 24.22
Qwen2.5-VL-32B 33.67 81.44 57.56 31.00
Qwen2.5-VL-7B 23.00 77.89 50.44 12.89
InternVL-2.5-38B 55.67 60.00 57.83 34.56
InternVL-2.5-8B 58.33 45.67 52.00 21.33
LLaVA-NeXT-7B 54.44 40.56 47.50 10.78
LLaVA-OneVision-7B 19.78 81.56 50.67 10.44
Table 2: Experimental results for the UI/UX design selection task on WiserUI-Bench. Best values are in bold; second-best are underlined. All metrics are reported as percentages (%).
Results
Table 2 summarizes model performance on the UI/UX design selection task. All AA scores barely clear the 50% random baseline for a binary choice. Moreover, they are inflated by strong position bias, especially in proprietary models and the Qwen2.5-VL family, which tend to pick the second image. When we examine CA, the order-invariant metric, performance drops back to near-random levels. Interestingly, Qwen2.5-VL-32B and InternVL-2.5-38B score slightly higher than proprietary models on CA, but still fall short of being adequate. Contrary to expectations, o1 shows the strongest bias and a lower CA than GPT-4o, suggesting that self-reasoning training of o1 does not transfer to this task. Smaller (7B/8B) models perform significantly worse than larger ones, confirming the scaling trend. Overall, current MLLMs struggle to consistently select the more effective UI/UX design in real-world UI pairs.
Figure 5: Qualitative results produced by o1, GPT-4o, and Qwen2.5-VL-32B on the two core tasks. (Left) Qwen2.5-VL-32B and GPT-4o consistently selected the correct and incorrect variants, respectively, whereas o1 showed position bias. (Right) o1 successfully captured all key rationales behind the winning design. Green/blue highlights mark alignment with expert rationales.
5.2 UI/UX Design Rationale Alignment
Metrics
To judge how well a model’s generated rationale explains the observed outcome, we compare it to a set of expert-curated key rationales per instance. We report two metrics: Rationale-wise accuracy, which checks model agreement with each expert rationale separately, and Instance-wise accuracy, which counts an instance as correct if the model matches any rationale. Alignment is assessed using a GPT-4o-based evaluator (Liu et al. 2023), validated against human judgments on 1,000 random samples (70.8% accuracy, Cohen’s kappa 0.42). We also report BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2020) in Supplementary.
Model Rationale-wise Instance-wise
o1 64.18 78.33
GPT-4o 50.15 66.67
Claude 3.5 Sonnet 63.16 78.67
Qwen2.5-VL-32B 60.23 73.00
Qwen2.5-VL-7B 43.71 61.67
InternVL-2.5-38B 45.32 59.33
InternVL-2.5-8B 35.09 51.67
LLaVA-NeXT-7B 11.99 21.67
LLaVA-OneVision-7B 21.78 33.33
Table 3: Results for the UI/UX design rationale alignment task on WiserUI-Bench, figures in percentages (%).
Results
Table 3 reports performance on the UI/UX design rationale alignment task. o1 and Claude 3.5 Sonnet stand out with expert-level explanations, showing strong post-hoc reasoning once the correct answer is given, even though they noticeably fell short on selection when the answer was not avaliable in Task 1. GPT-4o performs worse in this setting, while Qwen2.5-VL-32B remains solid and InternVL-2.5-38B drops. Qwen2.5-VL-7B shows relatively strong alignment, while other 7B/8B models continue to underperform. The consistent gap between instance-wise and rationale-wise scores indicates that while models can often align with part of the expert reasoning, they rarely capture it fully.
6 Analyses
6.1 Qualitative Analysis
We analyzed differences in model outputs in depth. The left part of Figure 5 illustrates an example from the UI/UX design selection task. Variant 1, the actual A/B test winner with larger product images that are more likely to capture users’ attention, was correctly chosen by Qwen2.5-VL-32B, while other models failed to do so. When the correct variant appeared first in the input, Qwen2.5-VL-32B provided appropriate reasoning, demonstrating a solid understanding of UI/UX. In contrast, both o1 and GPT-4o incorrectly attributed their choice to an increased number of items as the reason. Additionally, o1 exhibited position bias.
The right part of Figure 5 shows a case from the UI/UX design rationale alignment task. Given that the UI on the right was the A/B test-verified winner, experts identified the key reasons for its effectiveness as simplified background and enlargened demo button. both of which models must explicitly capture to provide an accurate explanation. o1 succeeds by directly highlighting these key points. GPT-4o correctly identifies the relevant UI elements, but provides an incorrect or vague explanation for their impact. Qwen2.5-VL-32B correctly notes the minimalist design, but stops at identifying the button, missing the key detail of its enlargement and resulting in an incomplete explanation.
6.2 Effect of Diverse Reasoning Strategies
For the first task, we additionally evaluate various reasoning strategies on GPT-4o and Claude 3.5 Sonnet to assess their effectiveness, including Contrastive Chain-of-Thought (CoCoT) (Zhang et al. 2024), Self-Refine (Madaan et al. 2023), Duty-Distinct Chain-of-Thought (DDCoT) (Zheng et al. 2023), and Multi-Agent Debate (MAD) (Liang et al. 2024; Du et al. 2023). For MAD, we stop and moderate the debate after the first (R1) and third (R3) rounds. As shown in Table 4, AA scores remain similar to or below the zero-shot setting, while CA scores often surpass it, especially for MAD (R1) across models, DDCoT on GPT-4o, and CoCoT on Claude 3.5 Sonnet. This gain is mainly due to reduced position bias, enabling more consistent decisions regardless of input order. MAD (R1) performs well by generating and moderating conflicting perspectives, which suits this human preference-driven task, benefiting from multi-perspective reasoning. However, excessive internal debate may hurt performance, as seen in the drop with MAD (R3).
@@ 6.2 Effect of Diverse Reasoning Strategies
Table 4: Results for the UI/UX design selection task across reasoning strategies on WiserUI-Bench; metrics match Section 5.1 and are averaged over three runs.
6.3 Rationale Alignment Across UX Aspects
We evaluated rationale-wise accuracy in the second task by each UX dimension, as truly understanding UI/UX requires capturing the full spectrum of cognitive processing—perception, memory, and action—reflected in our rationale categorization. As shown in Figure 6, o1 leads on Perception, indicating strong ability to reason about what users notice in a UI. Claude 3.5 Sonnet performs best on Action, suggesting it is more adept at predicting how users would interact with interfaces. Notably, model performance converge more on Memory than on the others. The persistent sub-70% ceiling across all dimensions underscores the need for deeper cognitive understanding of UX as a whole.
@@ 6.3 Rationale Alignment Across UX Aspects
Figure 6: Results for the UI/UX design rationale alignment task by each UX dimension on WiserUI-Bench.
7 Discussion and Conclusion
Toward Behavior-Aware UI Generation: Human Study on Synthetic Pairs
As an exploratory extension of our main study, we evaluated model predictions against a human study on synthetic UI pairs crafted to elicit specific user actions. We position this task and its results as a stepping stone toward automating behavior-aware UI generation.
Prior work on UI generation with large models largely targets matching natural-language or screenshot specifications (Laurençon, Tronchon, and Sanh 2024; Si et al. 2025; Li, Zhang, and Yang 2025) or ensuring basic functionality and visual polish (Wu et al. 2024b; Lu et al. 2025c), rarely examining whether the produced designs actually steer user behavior. Because influencing actions is the core purpose of UI design, generation must also reflect that goal.
Preference-based tuning methods such as Direct Preference Optimization (DPO) (Rafailov et al. 2023) may offer a plausible route. They enable a closed-loop pipeline without direct human supervision: (1) prompt with a behavioral goal, (2) generate candidates, (3) predict which better supports the goal, and (4) feed that signal back to refine the model (Yuan et al. 2024). For this to work, models must first reliably pick the more behavior-effective design not only on real-world interfaces, as in our main study, but also on synthetic pairs.
To test this prerequisite, we conducted a human study with 55 participants on GPT-4o-generated synthetic UI pairs (code and its screenshot) based on 10 behavioral objectives. Participants were shown the objective and chose the variant that better supports it, yielding a preference signal focused on behavior; the majority choice serves as its proxy. Models, in contrast, evaluated the same pairs without access to the goal, mirroring setup in our main study. See Supplementary for details.
Table 5 shows that both proprietary and open-source MLLMs, while most performing better on CA than in main study, still remain near random on AA in this synthetic evaluation, mirroring their limitations observed on WiserUI-Bench. Figure 7 provides a representative example. Going forward, consistently identifying the more effective design in such synthetic pairs will be crucial for behavior-aware UI generation, as it supplies the scalable preference signal needed for models’ closed-loop improvement. We encourage further work in this direction.
Figure 7: Example from the human study. Given the instruction with a behavioral goal, participants selected the better synthetic variant. InternVL2.5-38B failed to match the majority choice due to flawed judgment.
Model FA SA AA CA
GPT-4o 46.67 70.00 58.33 46.67
Claude 3.5 Sonnet 23.33 93.33 58.33 23.33
Qwen2.5-VL-32B 43.33 66.67 55.00 40.00
InternVL-2.5-38B 56.67 46.67 51.67 40.00
Table 5: Results of the models on identifying the behaviorally preferred synthetic UI variant in human study. Reported figures are averaged over three trials.
Conclusion
@@
We introduce a new task that evaluates multimodal understanding of how design drives user actions in the critical field of UI/UX. To support this, we present WiserUI-Bench, a behavior-focused benchmark uniquely built from diverse, real-world A/B-tested UI variants and expert-curated key rationales, providing a reliable basis for assessment. Our two tasks, selection and explanation, go beyond surface-level recognition, requiring models to reason about the behavioral implications of visual design, a cognitively challenging problem. Extensive experiments across both current proprietary and open-source MLLMs reveal they exhibit only shallow levels of nuanced visual reasoning. Achieving stronger performance on this significant task may, for example, require further training on vast amounts of behaviorally grounded UI/UX data that reflect real-world dynamics. We envision WiserUI-Bench and our work as a step toward multimodal systems capable of not only understanding but also generating behavior-aware interfaces, aligning with human cognition and preferences to optimize user experience. More broadly, we hope to catalyze future research toward deeper UI/UX and human-behavior literacy in AI.
A Benchmark for Multimodal UI/UX Design Understanding
Supplementary Material
Appendix A Overview
This supplementary material includes the following sections: (B) Limitations of our work, (C) Ethics Statements for UI/UX experts and human study participants, Detailed descriptions of (D) WiserUI-Bench, (E) experiments, and (F) human study, and concludes with (G) icon attribution.
Appendix B Limitation
@@
Given that user experience patterns vary across cultural contexts and societal norms, cultural biases inevitably exist in our benchmark (Karreman, Romeo, and Li 2016). However, accounting for such diversity is also part of the benchmark’s intent, making it a more realistic and challenging evaluation setting. Also, interactive UIs were not extensively explored in our work. Although the number of 300 UI image pairs may seem limited, publicly available real-world A/B test results are extremely rare, as companies rarely disclose them. Despite the dataset size, it covers a wide range of diverse UI changes, each accompanied by multiple expert-validated rationales.
Appendix C Ethics Statement
@@
To ensure expert-level annotation quality, we recruited three annotators with professional UI/UX backgrounds, all currently employed at global IT companies specializing UX and AI. They were clearly informed of the academic nature of the task and voluntarily participated in the study. Each annotator received monetary compensation equivalent to approximately \$225 USD per person, which reflects fair market payment for domain-specific labor. Payments were issued within 48 hours of annotation completion. All annotators were instructed to focus solely on task-relevant elements and were not exposed to or asked to label any sensitive, offensive, or triggering content.
The human study was also conducted with strict ethical considerations. To minimize biases, we employed a double-blind setup. All 55 participants voluntarily took part in the survey, fully informed that the study was for academic purposes. Additionally, the survey contained no explicit or problematic content, ensuring a safe and respectful research environment.
Appendix D WiserUI-Bench Details
D.1 Image Preprocess
@@
During benchmark construction process, we observed that some UI images from our data sources included manual indicators, such as circled regions, numbered markers, and arrows, highlighting specific UI elements related to reasoning points from A/B test results (see Figure J). While helpful for human interpretation, these indicators introduced artificial attention cues that could potentially bias model evaluation.
@@
D.2 Rationale Annotation
As shown in Figure L, three UI/UX experts are presented with a pair of UI images, one of which was designated as the winner based on A/B test results, along with a corresponding source website link. Experts were instructed to write key rationales explaining why the winning version supported user actions more effectively than the alternative. We engaged experts for annotation, as user behavior is often subconscious and shaped by implicit context, making evaluation by non-experts inherently unreliable. These rationales underwent multiple rounds of review to ensure that no important reason was omitted and no irrelevant reason was included.
Expert annotators were guided by the original sources’ textual description which often highlights key design intentions, and two types of theoretical heuristics selected for their robustness and relevance to user behavior analysis: (1) the Nielsen Norman 10 Usability Heuristics (Nielsen 1994), serving as broad conceptual criteria, and (2) 12 curated UX laws that are still widely applied in practice. These UX laws provide concrete, actionable guidance across diverse aspects of user experience, which can be categorized into three cognitive dimensions: Perception, Memory, and Action. This classification is inspired by the framework proposed by Norman (2007), which introduced the concept of Emotional Design. Emotional design focuses on creating products that evoke emotions and foster positive user experience by engaging users at three cognitive levels: visceral, behavioral, and reflective. The full list of UX laws and their corresponding classifications is presented in Table A.
D.3 Statistics
Figure K presents the distribution of WiserUI-Bench across page types, industry domains, device types, and types of UI element changes. For clarification on the page type classification, the homepage refers to the main entry point of a website, typically offering broad navigation and general content. In contrast, a landing page is usually a standalone page optimized for a specific user action, such as signing up or making a purchase. For clarification on the types of UI element changes, detailed definitions are provided in Table B.
Appendix E Experiment Details
E.1 Setup
Model Configuration

### v4

@@ 5 Experiments
We evaluate a diverse set of MLLMs, spanning both proprietary and open-source models. On the proprietary side, we include o1 (OpenAI et al., 2024b), GPT-4o (OpenAI et al., 2024a), Claude 3.5 Sonnet (Anthropic, 2024), all noted for their strong visual reasoning. For open-source models, we evaluate widely adopted models such as Qwen2.5-VL (7B, 32B) (Bai et al., 2025), InternVL-2.5 (8B, 38B) (Chen et al., 2025), LLaVA-NeXT-7B (Liu et al., 2024a) and LLaVA-OneVision-7B (Li et al., 2024a). Refer to Appendix B.1 for details.
@@ 5.1.1 Evaluation Metrics
We report model accuracies across input-order variations of each UI image pair, reflecting the multi-image input setting. Specifically, we report: (1) First Accuracy (FA), when the more effective UI/UX design is presented first in the input; (2) Second Accuracy (SA), when it is presented second; and (3) Average Accuracy (AA), the mean of FA and SA, reflecting overall correctness. Finally, we compute (4) Consistent Accuracy (CA), which checks whether the model selects the same, correct UI across both input orders of each pair. This metric isolates the model’s core visual reasoning ability by ensuring that its choices are based on content rather than position bias, a well-known issue in pairwise comparison settings (Wang et al., 2025b; Zhang et al., 2023). (Random guessing yields a CA of 25%: since there’s a 50% chance of being correct in each order.) All evaluations are conducted over three independent runs, and we report the averaged results. Experiment details, including the prompts, are provided in Appendix B.2.
@@
Table 3: Results of the UI/UX design selection task across reasoning strategies on WiserUI-Bench, figures in percentages (%).
5.1.2 Results
Overall Performance
Table 2 summarizes model performance on the UI/UX design selection task. Most AA scores hover only slightly above the 50% random baseline for a binary choice. Moreover, they are inflated by strong position bias, especially in proprietary models and the Qwen2.5-VL family, which tend to pick the second image. When we examine CA, the order-invariant metric, performance drops back to near-random levels. Interestingly, Qwen2.5-VL-32B and InternVL-2.5-38B score slightly higher than proprietary models on CA, but still fall short of being adequate. Contrary to expectations, o1 shows the strongest bias and a lower CA than GPT-4o, suggesting that self-reasoning of o1 does not transfer well to this task. Smaller (7B/8B) models perform significantly worse than larger ones, confirming the scaling trend. Overall, existing MLLMs struggle to consistently predict the more effective UI/UX design in real-world UI pairs.
Model Interpretation Instance-level
Recall Recall
o1 64.18 78.33
GPT-4o 50.15 66.67
Claude 3.5 Sonnet 63.16 78.67
Qwen2.5-VL-32B 60.23 73.00
Qwen2.5-VL-7B 43.71 61.67
InternVL-2.5-38B 45.32 59.33
InternVL-2.5-8B 35.09 51.67
LLaVA-NeXT-7B 11.99 21.67
LLaVA-OneVision-7B 21.78 33.33
Table 4: Results of the UI/UX design interpretation task on WiserUI-Bench, figures in percentages (%).
@@ 5.1.2 Results
Figure 5: Results of the UI/UX design interpretation task by each UX dimension on WiserUI-Bench.
Performance Across UI Element Types
The UI element-wise results in Table 2 show performance variations across change type categories, while preserving overall trends: GPT-4o and InternVL-2.5-38B generally achieve the highest AA and CA, respectively. Focusing on CA, Mixed cases, where changes span multiple functional UI element groups, perform best across most models, suggesting that aggregated multi-functional changes provide stronger behavioral cues than single-group modifications. Interactive Control (IC) and Content & State Display (CSD) exhibit comparable CA and rank below Mixed, indicating relatively moderate sensitivity to action-oriented (e.g., buttons) and content-level changes (e.g., images or text). By contrast, Container & Layout Structure (CLS) is the most challenging category, with sharply reduced CA across most models, reflecting limited sensitivity to structural layout changes, such as card or tile-level differences. Notably, Qwen2.5-VL-32B and InternVL-2.5-8B perform relatively better on CLS than other models, including GPT-4o.
Figure 6: Example responses from InternVL-2.5-38B on the UI/UX design selection task, showing model outputs under different input image orders for each case. Blue marks correct UI change perception, while red denotes incorrect or hallucinated perception. Purple indicates user behavior pattern-aligned reasoning, whereas misaligned reasoning in orange.
Figure 7: Example responses from o1, GPT-4o, and Qwen2.5-VL-32B on the UI/UX design interpretation task. Green and blue highlight points capturing the expert interpretations fully.
Effect of Reasoning Strategies
We additionally evaluate various reasoning strategies on GPT-4o and Claude 3.5 Sonnet to assess their effectiveness, including Contrastive Chain-of-Thought (CoCoT) (Zhang et al., 2024), Self-Refine (Madaan et al., 2023), Duty-Distinct Chain-of-Thought (DDCoT) (Zheng et al., 2023), and Multi-Agent Debate (MAD) (Liang et al., 2024; Du et al., 2023). For MAD, we stop and moderate the debate after the first (R1) and third (R3) rounds. As shown in Table 3, AA scores remain similar to or below the zero-shot setting, while CA scores often surpass it, especially for MAD (R1) across models, DDCoT on GPT-4o, and CoCoT on Claude 3.5 Sonnet. This gain is mainly due to reduced position bias, enabling more consistent decisions regardless of input order. MAD (R1) performs well by generating and moderating conflicting perspectives, which suits this human behavior pattern-driven task, benefiting from multi-perspective reasoning. However, excessive internal debate may hurt performance, as seen in the drop with MAD (R3).
5.2 UI/UX Design Interpretation
5.2.1 Evaluation Metrics
To judge how well a model’s generated interpretation explains the observed outcome, we compare it against expert-curated key interpretations for each instance. Specifically, we instruct the model to produce an unconstrained free-form explanation enumerating all relevant factors, and each expert interpretation is independently evaluated in a binary manner, whether it is fully captured by the explanation or not. We report two metrics: Interpretation Recall, measuring the proportion of expert interpretations captured, and Instance‑level Recall, which counts an instance as correct if at least one expert interpretation is captured. Binary inclusion judgments are made using a GPT-4o-based evaluator (Liu et al., 2023), validated against human judgments on 1,000 random samples (83.0% accuracy, Cohen’s kappa 0.66). The prompts used, along with the results of BLEU, ROUGE, METEOR, and BERTScore (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005; Zhang et al., 2020) are reported in the Appendix B.3.
5.2.2 Results
Overall Performance
Table 4 reports performance on the UI/UX design interpretation task. o1 and Claude 3.5 Sonnet exhibit relatively strong post-hoc interpretation capability when the more effective variant is provided, despite having noticeably underperformed on selection in Task 1. By contrast, GPT-4o and InternVL-2.5-38B perform poorly under this setting, while Qwen2.5-VL-32B performs reasonably well. Among smaller models, Qwen2.5-VL-7B shows relatively strong alignment, whereas other 7B/8B models underperform. The consistent gap between Interpretation and Instance-level Recall scores indicates that while models can often recover parts of the expert interpretations, they rarely capture them in full.
Performance Across UX Aspects
We analyze Interpretation Recall scores by UX dimension, examining which types of UX interpretations are correctly recovered, as a true understanding of UI/UX requires capturing the full range of cognitive aspects—perception, memory, and action. As shown in Figure 5, o1 leads on Perception, indicating stronger ability to explain which visual cues are likely noticed by users. Claude 3.5 Sonnet performs best on Action, suggesting greater sensitivity to interaction-oriented factors. Notably, model performance converges more on Memory than on the others. The persistent sub-70% ceiling across all dimensions underscores the need for deeper, more holistic modeling of user cognition in UI/UX understanding and interpretation.
6 Discussion
Case Study
For the UI/UX design selection task, we analyze the success and failure modes of the strong-performing model, InternVL-2.5-38B, as shown in Figure 6, focusing on cases where the model makes consistent decisions under both input image orders. In Case 1, the model correctly perceives the UI change as the presence of a CTA button and selects the more effective UI/UX, providing a reason that behaviorally aligns with the user behavior pattern implied by the A/B test outcome, hypothesizing that fewer CTAs better focus user attention. In Case 2, although the model correctly identifies differences such as pricing badges and CTA text, it fails to select the winning variant identified by the A/B test, instead offering misaligned reasoning that higher visual saliency negatively impacts user behavior by inducing commitment pressure. In Case 3, the model fails at perception level, hallucinating non-existent changes instead of recognizing the actual button color difference, which ultimately leads to an incorrect final decision.
Figure 7 shows a case from the UI/UX design interpretation task. The UI on the right was verified by A/B test results as the more effective design, with experts attributing its effectiveness to a simplified background and an enlarged demo button, along with their associated effects, factors that models must explicitly capture for an accurate interpretation. o1 succeeds by directly highlighting these key points. GPT-4o correctly identifies the relevant UI elements, but provides an incorrect or vague explanation for their impact. Qwen2.5-VL-32B correctly notes the minimalist design, but stops at identifying the button, missing the key detail of its enlargement and resulting in an incomplete explanation.
7 Conclusion
@@
We introduce a new task that evaluates multimodal understanding of how design drives user actions in the critical field of UI/UX. To support this, we present WiserUI-Bench, a behavior-focused benchmark uniquely built from diverse, real-world A/B‑tested UI variants and expert-curated key interpretations, providing a reliable basis for assessment. Our two tasks, selection and interpretation, go beyond surface-level recognition, requiring models to reason about the behavioral implications of visual design, a cognitively challenging problem. Extensive experiments across both proprietary and open-source MLLMs reveal they exhibit only shallow levels of behavior-aware visual reasoning. Achieving stronger performance on this significant task, particularly across the analyzed UI elements or UX dimensions, may require further training on large-scale, behaviorally grounded UI/UX data that reflects real-world user dynamics. We envision WiserUI-Bench and our work as a step toward multimodal systems capable of not only understanding but also generating visual interfaces grounded in user behavior contexts, aligning with human cognition to optimize user experience. More broadly, we hope to catalyze future research toward deeper UI/UX and human behavior literacy in AI.
8 Limitations
@@
Given that user experience patterns vary across cultural contexts and societal norms, cultural biases inevitably exist in our benchmark (Karreman et al., 2016). Also, interactive UIs were not extensively explored in our work. Although the number of 300 UI image pairs may seem limited, publicly available real-world A/B test results are extremely rare, as companies rarely disclose them. Moreover, independently conducting production-level UI experiments at a comparable level and scale was practically infeasible. However, as the first benchmark in this area, we focused on reliability, diversity, and annotation quality, covering a wide range of real-world UI changes and providing expert-curated key interpretations.
9 Ethical Considerations
@@
To ensure expert-level annotation quality, we recruited three annotators with professional UI/UX backgrounds who are currently employed at global IT companies specializing in UX and AI. We clearly informed them of the academic nature of the task, and they voluntarily participated in the study. We provided each annotator with monetary compensation equivalent to approximately \$225 USD, reflecting fair market rates for domain-specific labor, and issued payments within 48 hours of annotation completion. We instructed annotators to focus solely on task-relevant elements and ensured that they were not exposed to or asked to label any sensitive, offensive, or triggering content.
Our benchmark does not include any private, user-identifiable, or confidential information. All materials were collected from publicly released sources originally disclosed by the originating platforms or companies. Nevertheless, to avoid downstream commercial use beyond the original disclosure scope, we plan to release the benchmark under a non-commercial research license.
Appendix A WiserUI-Bench Details
A.1 Curation Details
Image Preprocess
@@ A.1 Curation Details
During benchmark construction process, we observed that some UI images from our data sources included manual indicators, such as circled regions, numbered markers, and arrows, highlighting specific UI elements related to reasoning points from A/B test results (see Figure 17). While helpful for human interpretation, these indicators introduced artificial attention cues that could potentially bias model evaluation.
@@
Interpretation Annotation
As shown in Figure 20, three UI/UX experts are presented with a pair of UI images, one designated as the winner based on A/B test results, along with a corresponding source website link. Experts were instructed to provide key interpretations explaining why the winning version supported user actions more effectively than the alternative. We engaged experts for annotation, as user behavior is often subconscious and shaped by implicit context, making evaluation by non-experts inherently unreliable. To distill salient behavioral drivers, we retained only interpretations independently identified by at least two experts, resulting in 684 final interpretations (562 with unanimous agreement and 122 with two-expert agreement) from an initial pool of 845 candidates.
Expert annotators were initially guided by the original sources’ textual description which often highlights key design intentions, as well as by a robust theoretical heuristic widely used in usability research, the Nielsen Norman 10 Usability Heuristics (Nielsen, 1994), which served as broad conceptual criteria. The resulting interpretations were then consolidated and mapped to 12 curated UX laws that remain widely applied in practice. These UX laws provide concrete, actionable guidance across diverse aspects of user experience, which can be categorized into three cognitive dimensions: Perception, Memory, and Action. This classification is inspired by the framework proposed by Norman (2007), which introduced the concept of Emotional Design. Emotional design focuses on creating products that evoke emotions and foster positive user experience by engaging users at three cognitive levels: visceral, behavioral, and reflective. The full list of UX laws and their corresponding classifications is presented in Table 5.
A.2 Statistics Details
Figure 18 presents the distribution of WiserUI-Bench across page types, industry domains, device types. For clarification on the page type classification, the homepage refers to the main entry point of a website, typically offering broad navigation and general content. In contrast, a landing page is usually a standalone page optimized for a specific user action, such as signing up or making a purchase.
Full distribution of UI change attribute types is shown in Figure 19. For clarification, element-level counts reflect the number of distinct UI components affected, while attribute-level counts capture the total number of modified attributes across all components. For example, an instance annotated as <Input Field: Padding, Color / Button: Color> is counted as 2 element changes and 3 attribute changes.
Appendix B Experiment Details
B.1 Setup
B.1.1 Model Configuration